OpenStudy (anonymous):

What is the difference between the right-hand limit and the left-hand limit of f(x) =|x − 3|/x − 3 as x approaches 3? no difference 1 unit 2 units 3 units

zepdrix (zepdrix):

From the left side of x=3,\[\Large\rm |x-3|=-(x-3)\]Plug that in for your left-sided limit. From the right side of x=3,\[\Large\rm |x-3|=x-3\]Plug that in for the right-sided limit :)

OpenStudy (anonymous):

i still dont understand. sorry i new to this

zepdrix (zepdrix):

\[\Large\rm |x-3|=\cases{x-3,&x>3\\ \Large\rm -(x-3), &x<3}\] \[\Large\rm \lim_{x\to3^-} \frac{|x-3|}{x-3}=\lim_{x\to3^-} \frac{-(x-3)}{x-3}\]The absolute function is negative when we're left of 3. Simplify that to get your left-sided limit.

zepdrix (zepdrix):

\[\Large\rm \lim_{x\to3^+}\frac{|x-3|}{x-3}=\lim_{x\to3^+}\frac{x-3}{x-3}\]|x-3| is positive when we're on the right side of 3. Simplify! :) What do you get for your two limits?

OpenStudy (jdoe0001):

\(\bf lim_{x\to 3}\ \cfrac{|x-3|}{x-3}\implies \begin{cases} lim_{x\to {\color{red}{ 3^-}}}\ \cfrac{-(x-3)}{x-3} \\ \quad \\ \bf lim_{x\to {\color{red}{ 3^+}}}\ \cfrac{+(x-3)}{x-3} \end{cases}\)

OpenStudy (anonymous):

i think its no difference right?

OpenStudy (anonymous):

@davidza6496 I thought @kirbykirby did this with you earlier?

OpenStudy (anonymous):

Try looking at the graph of \(\large y=\frac{|x-3|}{x-3} \) and compare the y-values both to the left and right of x=3.

OpenStudy (anonymous):

is the difference 2 i think i got

zepdrix (zepdrix):

So you got -1 from the left side, and 1 from the right side? distance of 2 between them? Yes good job! :)